(D^3 + 3D^2-D-3)y = e^2x+4


1
Expert's answer
2021-12-06T17:29:21-0500

Solution

For general solution of the homogeneous equation y’’’+ 3y’’- y’- 3y = 0 we need to solve characteristic equation

k3 + 3k2 – k – 3 = 0

(k + 3)(k2 – 1) = 0

(k + 3)(k – 1) (k + 1) = 0  => k1 = -3, k2 = 1, k3 = -1

So general solution of the homogeneous equation is y0(x) = Ae-3x + Bex + Ce-x , where A, B, C are arbitrary constants.

Let’s find partial solution of nonhomogeneous equation in the form y1(x) = Ee2x + F, where E, F are arbitrary constants. Substitution into equation gives

y1’ = 2Ee2x , y1’’ = 4Ee2x , y1’’’ = 8Ee2x

y1’’’+ 3y1’’- y1’- 3y1 = 8Ee2x + 12Ee2x – 2Ee2x – 3Ee2x – 3F = 15Ee2x  – 3F = e2x+4

So E = 1/15, F = -4/3

Therefore solution of the given equation is

y(x) = y0(x) + y1(x) = Ae-3x + Bex + Ce-x + e2x /15 – 4/3